# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v1

@@ Abstract
Principal component analysis (PCA) is a key tool in the field of data dimensionality reduction. Various methods have been proposed to extend PCA to the union of subspace (UoS) setting for clustering data that come from multiple subspaces like $K$-Subspaces (KSS). However, some applications involve heterogeneous data that vary in quality due to noise characteristics associated with each data sample. Heteroscedastic methods aim to deal with such mixed data quality. This paper develops a heteroscedastic-focused subspace clustering method, named ALPCAHUS, that can estimate the sample-wise noise variances and use this information to improve the estimate of the subspace bases associated with the low-rank structure of the data. This clustering algorithm builds on K-Subspaces (KSS) principles by extending the recently proposed heteroscedastic PCA method, named LR-ALPCAH, for clusters with heteroscedastic noise in the UoS setting. Simulations and real-data experiments show the effectiveness of accounting for data heteroscedasticity compared to existing clustering algorithms. Code available at https://github.com/javiersc1/ALPCAHUS.
@@ II Problem Formulation
Let $K$ denote the number of subspaces that is either known beforehand or estimated using other methods. Before describing the general union of subspace model, we consider the single subspace model when $K=1$.
@@ III-A1 Spectral Clustering
$h_{ij}\triangleq\left\{\begin{array}[]{ll}1/\sqrt{|\mathcal{A}_{j}|}&\text{ if}v_{i}\in\mathcal{A}_{j}\\ 0&\text{ otherwise}\end{array}\right.$
@@
where the notation s.t. means subject to some conditions. Trace minimization problems with semi-unitary constraints are well-studied in the literature, and it is easy to see that $\hat{\bm{H}}$ consists of the first $K$ eigenvectors of $\bm{L}$ assuming that $\lambda_{1}(\bm{L})\leq\ldots\leq\lambda_{N}(\bm{L})$ where $\lambda_{i}(\bm{L})$ denotes the $i$th eigenvalue of $\bm{L}$. Once $\hat{\bm{H}}$ is computed, the $K$-means algorithm is applied to $\hat{\bm{H}}^{T}$, treating each row of $\hat{\bm{H}}$ as the spectral embedding of the associated $v_{i}$ vertex. We now describe three subspace clustering methods that use this technique by first forming a weight matrix $\bm{W}$ and then applying the spectral clustering method.
@@ III-A2 Self-Expressive Methods
where $\|\cdot\|_{1,1}=\|\text{vec}(\cdot)\|_{1}$ is the vectorized 1-norm and $\left\|\cdot\right\|_{\mathrm{F}}$ is the Frobenius norm of a matrix. Observe that (8) approximates each data sample as a sparse linear combination of other data points to form $\bm{P}$. Let $\text{abs}(\bm{P})$ represent the element-wise absolute value of matrix $\bm{P}$. Then, spectral clustering is performed on $\bm{W}=\frac{1}{2}(\text{abs}(\bm{P}^{T})+\text{abs}(\bm{P}))$ by applying the $K$-means method to the spectral embedding of the affinity matrix $\bm{W}$. Ideally, (8) would select the high quality data to represent worse samples in $\bm{P}$ and this information would be retained in $\bm{W}$. However, this data quality awareness condition is not guaranteed in self-expressive methods to our knowledge. In our experiments, Fig. 2(f) and Fig. 6 show that there must be an issue constructing an ideal $\bm{P}$ due to poor clustering performance in heteroscedastic conditions.
@@ III-B Single Subspace Heteroscedastic PCA Method
Before extending (11) to the heteroscedastic setting, we review how LR-ALPCAH [17] solves for a single subspace ($K=1$) in the heteroscedastic setting using the model described in (1). For the measurement model $\bm{y}_{i}\sim\mathcal{N}(\bm{x}_{i},\nu_{i}\bm{I})$ in (1), the probability density function for a single data sample $\bm{y}_{i}$ is
@@ IV Proposed Subspace Clustering Method
For notational simplicity, let $\bm{Y}_{k}\triangleq\bm{Y}_{\mathcal{C}_{k}}\in\mathbb{R}^{D\times N_{k}}$ denote the submatrix of $\bm{Y}$ having columns corresponding to data samples that are estimated to belong in the $k$th subspaces, i.e., $\bm{Y}_{k}=\bm{Y}_{\mathcal{C}_{k}}=\text{matrix}(\{\bm{y}_{i}\,:\,c_{i}=k\})$. We apply this notation similarly to other matrices such as $\bm{\Pi}_{k}\triangleq\bm{\Pi}_{\mathcal{C}_{k}}=\text{diag}(\{\nu_{i}\,:\,c_{i}=k\})$. For the union of subspace measurement model (2), we generalize LR-ALPCAH (14) with the following optimization problem
@@
where $\mathcal{C},\bm{\Pi},\mathcal{L},\mathcal{R}$ denote the sets of estimated clusters, noise variances, and factorized matrices respectively for each cluster $k=1,\ldots,K$. Specifically $\mathcal{L}\triangleq\{\bm{L}_{1},\ldots,\bm{L}_{K}\},\ \mathcal{R}\triangleq\{\bm{R}_{1},\ldots,\bm{R}_{K}\}.$ Our algorithm for solving (15) is called ALPCAHUS (ALPCAH for Union of Subspaces). Similar to KSS, we alternate between updating subspace bases given cluster estimates and updating cluster estimates by projection given subspace bases estimates. We solve (15) via alternating minimization between basis estimation $\forall k\in\{1,\ldots,K\}$ solve
@@
and data sample reassignment $\forall i\in\{1,\ldots,N\}$ solve
@@
$z_{k}(j)\triangleq\max(\alpha,|\mathcal{C}_{k}|^{-1}\|(\bm{Y}_{k}-\bm{L}^{(t_{1})}_{k}\bm{R}^{(t_{1})^{T}}_{k})\bm{e}_{j}\|_{2}^{2})$ (20)
@@
$=\text{diag}(z_{k}(1),\ldots,z_{k}(N)).$ (21)
Because $\bm{\Pi}_{k}$ is a diagonal matrix, the $\bm{\Pi}_{k}\succeq\alpha\bm{I}$ condition in (16) leads to a projection to the positive orthant, i.e., the $\max(\alpha,\cdot)$ condition in (20). This condition is necessary to ensure convergence as proven in [17, Thm. 2] and used in this work to further prove ALPCAHUS convergence in Thm. 1 using some $\alpha\in\mathbb{R}_{+}$.
The solution to (17) involves the following procedure. Given a data sample $\bm{y}_{i}$, find the lowest subspace projection residual out of all subspaces and assign it to $c_{i}^{(t_{2}+1)}$ at the $(t_{2}+1)$ iteration. In the event of a tie, meaning there is more than one subspace that has equal residual, copy the past label $c_{i}^{(t_{2})}$ to $c_{i}^{(t_{2}+1)}$. In (17), $\mathcal{S}_{J_{i}}$ refers to the set of possible $k$ that achieves the lowest cost $J_{i}(k)$ in the event of a tie, i.e., $J_{i}(k_{1})=J_{i}(k_{2})\text{ for }k_{1}\neq k_{2}$. By doing this, cycling between cluster labels is prevented for all data samples. This cluster reassignment criteria will play a big role in ensuring convergence as proven in Thm. 1.
To further improve clustering performance, we leverage consensus clustering over many trials. Initially, we tried using this approach with only one trial as seen in Fig. 2(b) and ALPCAHUS ($B=1$) result in Fig. 6(a). However, since $K$-subspaces in general is sensitive to initialization, we found great success in using a consensus approach with more than one trials as shown in Fig. 2(d) and ALPCAHUS ($B=16$) result in Fig. 6(a).
@@ IV-A Ensemble Extension for ALPCAHUS
We now present the ensemble algorithm with base clustering parameter $B$ to combine multiple trials of finding $\mathcal{C}$ in (15). Any $B>1$ leverages consensus clustering by forming an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ where
@@ IV-B Rank Estimation
This is done for each estimated cluster $\bm{Y}_{k}$ over $R$ random trials where $1\leq R\ll T_{2}=\text{ALPCAHUS maximum total iterations}$. Here $\sigma_{d+1}(\bm{Y}_{k})$ denotes the $d+1$th singular value of $\bm{Y}_{k}$. We repeat this process for all cluster subsets $\{\bm{Y}_{1},\ldots,\bm{Y}_{K}\}$ after the cluster reassignment update in (17). To reduce computation, we perform this dimension estimate sparingly every $10\%$ of ALPCAHUS iterations.
@@ IV-C Cluster Initialization
$\bm{W}_{ij}=1\text{ if }|\langle\bm{y}_{i},\bm{y}_{j}\rangle|\geq q\text{ and }i\neq j$ (28)
given a thresholding parameter $q>0$ and zero otherwise. Then, the cluster assignments $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$ are calculated by applying the spectral clustering method on $W$. Recall that $\bm{y}_{i}=\bm{x}_{i}+\bm{\epsilon}_{i}$. Upon closer analysis, this metric in expectation, ignoring absolute value, gives
@@ IV-D ALPCAHUS Convergence
ALPCAHUS with a single cluster ($K=1$) has a sequence of cost function values that provably converges to local minima, since the cost function and algorithm simplifies to LR-ALPCAH (14) that has convergence guarantees shown in [17, Thm. 2]. In the multi-cluster setting ($K>1$), we show here that one can guarantee convergence of the sequence of cost function values in (15) by ensuring two things: 1) that the noise variances are kept above some small threshold value and 2) that the cluster assignment criteria makes assignment changes only if the cost function strictly decreases. The argument in Thm. 1 below is a natural extension of that in the original k-subspaces paper [25, Thm. 7]. The Appendix provides an experimental result, Fig. 9, that corroborates this theorem.
@@
Consider the base non-ensemble version of the ALPCAHUS algorithm. Assume a stopping criteria where we continue the loop over $t_{2}$ in the algorithm until there is an iterate where the cost function does not decrease. Assume a noise variance lower bound parameter $\alpha\in\mathbb{R}>0$ that lower bounds all $\nu_{i}$ and a cluster assignment criteria that accepts changes only if there is a cluster reassignments of points that strictly decrease the cost function, as expressed in (17). Then, the ALPCAHUS algorithm generates a sequence of cost function values that converges to a fixed point.
@@
To prove the cost function values converge, we need to assess whether each step decreases the cost function, and whether the algorithm terminates at a fixed point where the variables are no longer updated. Each step consists of two sub-steps: first, the subspace estimation sub-step where we perform LR-ALPCAH, and second, the clustering sub-step where we project each data vector onto the estimated subspaces and cluster with respect to smallest residual. For the subspace estimation sub-step, with cluster assignments fixed, we know that LR-ALPCAH has a sequence of cost function values that will converge according to guarantees given in [17, Thm. 2]. For the clustering sub-step, given our stopping criteria, once the clusters do not change our algorithm has a sequence of cost function values that reached a fixed point. That leaves only the case where the clusters do change, which we will focus on in the remainder of the proof. Let $\bm{Y}_{k}^{(t)},\bm{L}_{k}^{(t)},\bm{R}_{k}^{(t)},\bm{\Pi}_{k}^{(t)}$ be the optimization variables that produce the cost
@@
It is easy to see that $\bm{X}_{k}=\bm{L}_{k}\bm{R}_{k}^{T}=\bm{U}_{k}\bm{U}_{k}^{T}\bm{Y}_{k}$ since the projection of $\bm{Y}_{k}$ onto $\mathcal{R}(\bm{U}_{k})$ is $\bm{X}_{k}$. Then, cost can be expressed as
@@ V-A1 Experimental Setup
We generated a synthetic dataset consisting of $K=2$ clusters each of dimension $d=3$ derived from random subspaces in $D=100$ dimensional ambient space. Each cluster consisted of two data groups with group 1 containing $N_{1}=6$ samples, to explore the data constrained regime, with noise variance $\nu_{1}=0.1$ per cluster. For group 2, we varied the $N_{2}$ samples and noise variance $\nu_{2}$. We performed cross validation for hyperparameters in any algorithms that require it with a separate training set. We used the estimated clusters from each clustering algorithm and the known clusters to calculate clustering error using the Hungarian algorithm [35] to deal with the problem of label permutations. More concretely, clustering error is defined as
@@
We tried different values of noise variances and data samples and computed average clustering error out of 100 trials. Each trial had different noise, basis coefficients, and subspace basis realizations for the clusters. We compared various subspace clustering algorithms such as $K$-Subspaces (KSS) [14], Ensemble $K$-Subspaces (EKSS) [16], Subspace Clustering via Thresholding (TSC) [36], and Doubly Stochastic Sparse Subspace Clustering (ADSSC) [37]. Applying general non-subspace clustering methods such as $K$-means [38] resulted in poor clustering error so results of these methods are generally not shown.

### v2

@@ Abstract
Principal component analysis (PCA) is a key tool in the field of data dimensionality reduction. Various methods have been proposed to extend PCA to the union of subspace (UoS) setting for clustering data that comes from multiple subspaces like $K$-Subspaces (KSS). However, some applications involve heterogeneous data that vary in quality due to noise characteristics associated with each data sample. Heteroscedastic methods aim to deal with such mixed data quality. This paper develops a heteroscedastic-focused subspace clustering method, named ALPCAHUS, that can estimate the sample-wise noise variances and use this information to improve the estimate of the subspace bases associated with the low-rank structure of the data. This clustering algorithm builds on K-Subspaces (KSS) principles by extending the recently proposed heteroscedastic PCA method, named LR-ALPCAH, for clusters with heteroscedastic noise in the UoS setting. Simulations and real-data experiments show the effectiveness of accounting for data heteroscedasticity compared to existing clustering algorithms. Code available at https://github.com/javiersc1/ALPCAHUS.
@@ II Problem Formulation
Let $K$ denote the number of subspaces that is either known beforehand or estimated using other methods. Before describing the general union of subspace model, the single subspace model ($K=1$) is introduced.
@@ III-A1 Spectral Clustering
$h_{ij}\triangleq\left\{\begin{array}[]{ll}1/\sqrt{|\mathcal{A}_{j}|}&\text{ if}v_{i}\in\mathcal{A}_{j}\\ 0&\text{ otherwise}\end{array}\right\}$
@@
where the notation s.t. means subject to some conditions. Trace minimization problems with semi-unitary constraints are well-studied in the literature, and it is easy to see that $\hat{\bm{H}}$ consists of the first $K$ eigenvectors of $\bm{L}$ assuming that $\lambda_{1}(\bm{L})\leq\ldots\leq\lambda_{N}(\bm{L})$ where $\lambda_{i}(\bm{L})$ denotes the $i$th eigenvalue of $\bm{L}$. Once $\hat{\bm{H}}$ is computed, the $K$-means algorithm is applied to $\hat{\bm{H}}^{T}$, treating each row of $\hat{\bm{H}}$ as the spectral embedding of the associated $v_{i}$ vertex. The next section, Sec. III-A2, now describes some subspace clustering methods that use this technique by first forming a weight matrix $\bm{W}$ and then applying the spectral clustering method.
@@ III-A2 Self-Expressive Methods
where $\|\cdot\|_{1,1}=\|\text{vec}(\cdot)\|_{1}$ is the vectorized 1-norm and $\left\|\cdot\right\|_{\mathrm{F}}$ is the Frobenius norm of a matrix. Observe that (8) approximates each data sample as a sparse linear combination of other data points to form $\bm{P}$. Let $\text{abs}(\bm{P})$ represent the element-wise absolute value of matrix $\bm{P}$. Then, spectral clustering is performed on $\bm{W}=\frac{1}{2}(\text{abs}(\bm{P}^{T})+\text{abs}(\bm{P}))$ by applying the $K$-means method to the spectral embedding of the affinity matrix $\bm{W}$. Ideally, (8) would select the high quality data to represent worse samples in $\bm{P}$ and this information would be retained in $\bm{W}$. However, this data quality awareness condition is not guaranteed in self-expressive methods to our knowledge. In our experiments, Fig. 2(f) and Fig. 6 show that there must be an issue constructing an ideal $\bm{P}$ due to worse clustering performance in heteroscedastic conditions.
@@ III-B Single Subspace Heteroscedastic PCA Method
Before extending (11) to the heteroscedastic setting, a review of how LR-ALPCAH [17] solves for a single subspace ($K=1$) in the heteroscedastic setting using the model described in (1) is necessary. For the measurement model $\bm{y}_{i}\sim\mathcal{N}(\bm{x}_{i},\nu_{i}\bm{I})$ in (1), the probability density function for a single data sample $\bm{y}_{i}$ is
@@ IV Proposed Subspace Clustering Method
For notational simplicity, let $\bm{Y}_{k}\triangleq\bm{Y}_{\mathcal{C}_{k}}\in\mathbb{R}^{D\times N_{k}}$ denote the submatrix of $\bm{Y}$ having columns corresponding to data samples that are estimated to belong in the $k$th subspaces, i.e., $\bm{Y}_{k}\triangleq\bm{Y}_{\mathcal{C}_{k}}\triangleq\text{matrix}(\{\bm{y}_{i}\,:\,c_{i}=k\})$. This notation applies similarly to other matrices such as $\bm{\Pi}_{k}\triangleq\bm{\Pi}_{\mathcal{C}_{k}}\triangleq\text{diag}(\{\nu_{i}\,:\,c_{i}=k\})$. For the union of subspace measurement model (2), we generalize LR-ALPCAH (14) with the following optimization problem
@@
where $\mathcal{C},\bm{\Pi},\mathcal{L},\mathcal{R}$ denote the sets of estimated clusters, noise variances, and factorized matrices respectively for each cluster $k=1,\ldots,K$. Specifically $\mathcal{L}\triangleq\{\bm{L}_{1},\ldots,\bm{L}_{K}\}$ and $\mathcal{R}\triangleq\{\bm{R}_{1},\ldots,\bm{R}_{K}\}$. Our algorithm for solving (15) is called ALPCAHUS (ALPCAH for Union of Subspaces). Similar to KSS, ALPCAHUS alternates between updating subspace bases given cluster estimates and updating cluster estimates by projection given subspace bases estimates. We solve (15) with the alternating minimization technique. The method alternates between basis estimation, $\forall k\in\{1,\ldots,K\}$ solve
@@
and data sample reassignment, $\forall i\in\{1,\ldots,N\}$ solve
@@
$p_{k}(j)\triangleq\max(\alpha,|\mathcal{C}_{k}|^{-1}\|(\bm{Y}_{k}-\bm{L}^{(t_{1})}_{k}\bm{R}^{(t_{1})^{T}}_{k})\bm{e}_{j}\|_{2}^{2})$ (20)
@@
$=\text{diag}(p_{k}(1),\ldots,p_{k}(|\mathcal{C}_{k}|)).$ (21)
Because $\bm{\Pi}_{k}$ is a diagonal matrix, the $\bm{\Pi}_{k}\succeq\alpha\bm{I}$ majorization condition in (16) being equivalent to $\forall i,\lambda_{i}(\bm{\Pi}_{k})\geq\alpha$ implies that $\forall i,\nu_{i}>\alpha$. This leads to a projection to the positive orthant, i.e., the $\max(\alpha,\cdot)$ condition in (20). This condition is necessary to ensure convergence as proven in [17, Thm. 2] and used in this work to further prove ALPCAHUS convergence in Thm. 1 using some $\alpha\in\mathbb{R}_{+}$.
The solution to (17) involves the following procedure. Given a data sample $\bm{y}_{i}$, find the lowest subspace projection residual out of all subspaces and assign it to $c_{i}^{(t_{2}+1)}$ at the $(t_{2}+1)$ iteration. In the event of a tie, meaning there is more than one subspace that has equal residual, copy the past label $c_{i}^{(t_{2})}$ to $c_{i}^{(t_{2}+1)}$. In (17), $\mathcal{S}_{J_{i}}$ refers to the set of possible $k$ that achieves the lowest cost $J_{i}(k)$, i.e., $J_{i}(k_{1})=J_{i}(k_{2})\text{ for }k_{1}\neq k_{2}$. By doing this, cycling between cluster labels is prevented for all data samples. This cluster reassignment criteria will play a big role in ensuring convergence as proven in Thm. 1.
To further improve clustering performance, consensus clustering can be leveraged over many trials. Initially, we tried using this approach with only one trial as seen in Fig. 2(b) and ALPCAHUS ($B=1$) result in Fig. 6(a). However, since $K$-subspaces in general is sensitive to initialization, during experimentation there was great success in using a consensus approach with more than one trial as shown in Fig. 2(d) and ALPCAHUS ($B=16$) result in Fig. 6(a).
@@ IV-A Ensemble Extension for ALPCAHUS
This section presents the ensemble algorithm with base clustering parameter $B$ to combine multiple trials for finding $\mathcal{C}$ in (15). Any $B>1$ leverages consensus clustering by forming an affinity matrix $\bm{W}\in\mathbb{R}^{N\times N}$ where
@@ IV-B Rank Estimation
This is done for each estimated cluster $\bm{Y}_{k}$ over $R$ random trials where $1\leq R\ll T_{2}=\text{ALPCAHUS maximum total iterations}$. Here $\sigma_{d+1}(\bm{Y}_{k})$ denotes the $d+1$th singular value of $\bm{Y}_{k}$. This process is repeated for all cluster subsets $\{\bm{Y}_{1},\ldots,\bm{Y}_{K}\}$ after the cluster reassignment update in (17). To reduce computation, this dimension estimation is performed sparingly every $10\%$ of ALPCAHUS iterations.
@@ IV-C Cluster Initialization
$\bm{W}_{ij}=1\text{ if }|\langle\bm{y}_{i},\bm{y}_{j}\rangle|\geq\tau\text{ and }i\neq j$ (28)
given a thresholding parameter $\tau>0$ and zero otherwise. Then, the cluster assignments $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$ are calculated by applying the spectral clustering method on $\bm{W}$. Recall that $\bm{y}_{i}=\bm{x}_{i}+\bm{\epsilon}_{i}$. Upon closer analysis, this metric in expectation, ignoring absolute value, gives
@@ IV-D ALPCAHUS Convergence
ALPCAHUS with a single cluster ($K=1$) has a sequence of cost function values that provably converges to local minima, since the cost function and algorithm simplifies to LR-ALPCAH (14) that has convergence guarantees shown in [17, Thm. 2]. In the multi-cluster setting ($K>1$), Thm. 1 shows that one can guarantee convergence of the sequence of cost function values in (15) by ensuring two things: that the noise variances are kept above some small threshold value, and that the cluster assignment criteria makes assignment changes only if the cost function strictly decreases. The argument in Thm. 1 below is a natural extension of that in the original k-subspaces paper [25, Thm. 7]. The Appendix provides an experimental result, Fig. 9, that corroborates this theorem.
@@
Consider the non-ensemble version of the ALPCAHUS algorithm. Assume a stopping criteria where the loop over $t_{2}$ in the algorithm continues until there is an iterate where the cost function does not decrease. Assume a noise variance threshold parameter $\alpha\in\mathbb{R}>0$ that lower bounds all $\nu_{i}$ and a cluster assignment criteria that accepts changes only if there is a cluster reassignments of points that strictly decrease the cost function, as expressed in (17). Then, the ALPCAHUS algorithm generates a sequence of cost function values that converges to a fixed point.
@@
To prove that the sequence of cost function values converges, one needs to assess whether each step decreases the cost function, and whether the algorithm terminates at a fixed point where the variables are no longer updated. Each step consists of two sub-steps: first, the subspace estimation sub-step where LR-ALPCAH is used, and second, the clustering sub-step where projection onto the estimated subspaces is done to cluster with respect to the smallest residual. For the subspace estimation sub-step, with cluster assignments fixed, it is known that LR-ALPCAH has a sequence of cost function values that will converge according to guarantees given in [17, Thm. 2]. For the clustering sub-step, given our stopping criteria, once the clusters do not change, the algorithm has a sequence of cost function values that reaches a fixed point. That leaves only the case where the clusters do change, which will be the focus in the remainder of the proof. Let $\bm{Y}_{k}^{(t)},\bm{L}_{k}^{(t)},\bm{R}_{k}^{(t)},\bm{\Pi}_{k}^{(t)}$ be the optimization variables that produce the cost
@@
It is easy to see that $\bm{X}_{k}=\bm{L}_{k}\bm{R}_{k}^{T}=\bm{U}_{k}\bm{U}_{k}^{T}\bm{Y}_{k}$ since the projection of $\bm{Y}_{k}$ onto $\mathcal{R}(\bm{U}_{k})$ is $\bm{X}_{k}$. The cost can be expressed as
@@ V-A1 Experimental Setup
A synthetic dataset is generated consisting of $K=2$ clusters each of dimension $d=3$ derived from random subspaces in $D=100$ dimensional ambient space. Each cluster consisted of two data groups with group 1 containing $N_{1}=6$ samples, to explore the data constrained regime, with noise variance $\nu_{1}=0.1$ per cluster. For group 2, the $N_{2}$ samples and noise variance $\nu_{2}$ are varied. Cross validation is done for hyperparameters in any algorithms that require it with a separate training set. Both the estimated clusters from each clustering algorithm and the known clusters are necessary to calculate clustering error. Additionally, the problem of label permutations is dealt by using the Hungarian algorithm [35]. More concretely, clustering error is defined as
@@
The average clustering error is computed out of 100 trials where each trial had different noise, basis coefficients, and subspace basis realizations for the clusters. Various subspace clustering algorithms are compared such as $K$-Subspaces (KSS) [14], Ensemble $K$-Subspaces (EKSS) [16], Subspace Clustering via Thresholding (TSC) [36], and Doubly Stochastic Sparse Subspace Clustering (ADSSC) [37]. Applying general non-subspace clustering methods such as $K$-means [38] resulted in poor clustering error so results of these methods are generally not shown.